COMPUTING LOGARITHMIC CAPACITY WITH LINEAR PROGRAMMING

JÉRÉMIE ROSTAND

Abstract. We construct an algorithm for estimating the capacity of nice compact subsets of the plane. Our method is based on a relation between capacity and Green’s function together with a uniform harmonic approximation theorem. We use the simple least-square technique and, what is most important, we can obtain a bound on the error made in the estimated capacity.